# DenseDPO: Fine-Grained Temporal Preference Optimization for Video Diffusion Models

### v1

@@ 4.3 DPO with VLM Labels
VLM labelers. As discussed in Sec. 3.3, we aim to evaluate the effectiveness of pre-trained VLMs in video preference learning. We take VLMs designed specifically for the video quality assessment task, VideoScore [22], LiFT [72], VideoReward [45], and VisionReward [78]. They have been fine-tuned on large-scale human preference labels. We also utilize the state-of-the-art visual reasoning model, GPT o3 [1], to explore the limits of models without task-specific training. Finally, we design GPT o3 Segment that partitions long videos into short segments to process separately, and aggregates results via majority voting. Please refer to Appendix A.6 for more implementation details.
@@
Results. Tab. 3(a) presents the results of preference prediction. With task-specific fine-tuning, state-of-the-art video reward models outperform the advanced GPT o3 in assessing short clips. Yet, their performance on long videos degrades drastically. Thanks to the temporal alignment of video pairs, we can run GPT to compare short segments individually, and then aggregate the results. This leads to higher accuracy in long video preference prediction, without any preference data fine-tuning.
We further show the DPO alignment results in Tab. 3(b). Due to a higher preference accuracy, DenseDPO with GPT o3 Segment labels outperforms StructuralDPO with binary preferences from other VLMs. It even matches DenseDPO trained with human labels on text alignment, visual quality, and dynamic degree. Yet, please note that GPT label has 5.5$\times$ more videos than human label.
@@ 4.4 Ablation Study
Label bias. We first study if there is a systematic bias in the annotation pipeline, e.g., labelers may produce higher quality preference labels on short segments than long videos. To verify it, we aggregate labels of all segments within a video to obtain its binary preference label via majority voting, and then train StructuralDPO on these labels. Tab. 4 shows that it achieves similar results compared to StructuralDPO trained on globally labeled preferences. This proves that DenseDPO’s superior performance comes from segment-level preference supervision instead of labeler bias.
Label quality. We study how label noise affects DenseDPO performance. As shown in Tab. 4, we randomly flip 20% and 40% winning or losing labels. This results in a clear drop in all the metrics.
Label quantity. We ablate different amounts of segment preference labels used in DenseDPO. Tab. 4 shows that scaling to 2x labels leads to the best results across all metrics. Interestingly, results trained with 0.5$\times$ labels are clearly better than results trained with 40% labels flipped. This may indicate that the quality of labels has a larger impact than the quantity of labels.
@@ 5 Conclusion
We present DenseDPO, an improved preference optimization framework for video generation. We address two critical aspects of video DPO—comparison data curation and preference labeling. Our guided video generation mechanism and fine-grained segment-level labels lead to better post-training results compared to DPO baselines. We discuss our limitations in Appendix D.
@@ A.2 Baselines
Pre-trained model. Our base text-to-video model is a latent Diffusion Transformer framework [53]. It leverages a MAGVIT-v2 [82] as the autoencoder and a stack of 32 DiT blocks as the denoiser $\mathbf{G}_{\bm{\theta}}$. The autoencoder is similar to the one in CogVideoX [81], which downsamples the spatial dimensions by 8$\times$ and the temporal dimension by 4$\times$. Each DiT block is similar to the one in Open-Sora [36], which consists of a 3D self-attention layer running on all video tokens, a cross-attention layer between video tokens and T5 text embeddings [60] of the input prompt, and an MLP.
@@
SFT fine-tunes the pre-trained model on the 55k high-quality dataset described in Sec. A.1 for 5k iterations. We did not observe a clear difference between full-model and LoRA [28] fine-tuning, and thus choose LoRA fine-tuning as it is more efficient.
@@
DPO training hyper-parameters. Both VanillaDPO and StructuralDPO adopt the Flow-DPO loss with a constant $\beta$, as prior work shows it achieves the best performance [45]. Following prior works [78, 45], we use $\beta=500$ and apply LoRA [28] with rank 128 to fine-tune the video model. We train with the AdamW optimizer [48] and a global batch size of 256 for 1k steps. The peak learning rate is set to $1\times 10^{-5}$ and it is linearly warmed up from 00 in the first 250 steps. A gradient clipping of 1.0 is applied to stabilize training. We implement all models using PyTorch [52] and conduct training on 64 NVIDIA A100 GPUs, which takes around 16 hours.
Inference hyper-parameters. We use the rectified flow sampler [47] with 40 sampling steps and a classifier-free guidance [23] scale of $8$ to generate horizontal videos of $512\times 288$ resolution. A timestep shifting [17] of $5.66$ is applied to further improve results.
@@ A.4 Evaluation Datasets
Since our main focus is improving the temporal quality of the pre-trained video model and VanillaDPO, we evaluate on two motion-rich prompt sets.
@@
MotionBench collects more diverse prompts from existing prompt sets [56, 35, 72]. We run GPT-4o to select prompts with dynamic actions as described in Sec. A.1, resulting in 419 prompts. This prompt set contains more diverse scenes, subjects, and action types.
@@ B.1 Analyzing the Learning Signal
As discussed in Sec. 3.2, guided sampling is attractive since it fixes the structure in the preference pair. This neutralizes the motion bias between videos and focuses the comparison on visual artifacts. However, StructuralDPO with guided sampling achieves inferior results. Analyzing its learning signal reveals that it can paradoxically push the model to “unlearn” the real data distribution. Intuitively, this happens because the learning signal from a losing sample typically dominates over the winning one in those regions of a video, which correspond to the real data distribution. In this section, we investigate this phenomenon and discuss potential remedies to the issue.
@@
$(\displaystyle\mathcal{L}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},\,t\sim\mathcal{U}(0,T),\,\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}\mid\bm{x}_{0}^{w}),\,\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}\mid\bm{x}_{0}^{l})}\Big{[}\log\sigma\Big{(}-\beta T\big{(}$8)
@@
$(\displaystyle\mathcal{L}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},\,t\sim\mathcal{U}(0,T),\,\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}\mid\bm{x}_{0}^{w}),\,\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}\mid\bm{x}_{0}^{l})}\Big{[}\log\sigma\Big{(}-\beta T\omega(t)\big{(}$11)
@@
In guided sampling (see Algo. 2), we generate a winning/losing sample $\bm{x}^{*}_{n}=(1-\eta)\bm{x}_{0}+\eta\bm{\epsilon}^{*}$ using a real video $\bm{x}_{0}$. This makes them carry similar structure, which means that the winning and losing samples share many pixels. Moreover, these shared pixels normally correspond to the ground-truth data distribution. Let’s split the pixels into two sets, $\mathcal{I}_{\text{same}}$ and $\mathcal{I}_{\text{unique}}$:
$\mathcal{I}_{\text{same}}(\bm{x}^{w},\bm{x}^{l})$ =$=\{p\mid\bm{x}^{w}_{0}[p]\approx\bm{x}^{l}_{0}[p]\approx\bm{x}_{0}[p]\},$ (18)
$\mathcal{I}_{\text{unique}}(\bm{x}^{w}_{0},\bm{x}^{l}_{0})$ =$=\{p\mid p\not\in\mathcal{I}_{\text{same}}\},$ (19)
where $p$ is a pixel location (e.g., typically, triplet $(i,j,k)$ denoting the frame, height, width indices). In this way, $\mathcal{I}_{\text{same}}$ stores pixel locations which remained intact during the forward diffusion process of guided sampling and subsequent denoising with $\mathbf{G}_{\bm{\theta}}$. Fig. 2 and Fig. 11 illustrate this: many pixels in the winning and losing samples are identical and correspond to the original ground-truth video. In this way, one can argue that $\mathcal{I}_{\text{same}}$ corresponds to the real data distribution.
@@ B.2 Empirical Verification
The resulting dataset comprises five image variants, progressively more corrupted, visualized in Fig. 6 (top). This setup allows separate measurement of losses in corrupted and uncorrupted regions. Subsequently, we randomly sample a timestep $t\sim\mathcal{U}[0,1]$, estimate velocities $\bm{v}_{\text{Flux}}(\bm{x}_{t}^{w},t)$ and $\bm{v}_{\text{Flux}}(\bm{x}_{t}^{l},t)$, and compute per-pixel losses (averaged over the 16 latent channels $C$):
@@
Loss computations use the same noise seed for the corrupted and uncorrupted images (as is usually done in practice). Fig. 6 (top) illustrates a representative example of corruption. Visualizations in Fig. 6 (bottom) clearly demonstrate the loss dominance of the losing samples in the uncorrupted regions. As we maximize the loss of the losing sample and minimize that of the winning sample, this results in the model increasing loss in uncorrupted image regions when trained with DPO to mitigate blurring artifacts.
Fig. 7 provides a quantitative analysis, demonstrating how increasing artifact severity intensifies the negative learning signal from losing samples in uncorrupted image areas.
@@ Appendix C Additional Experimental Results
Figure 8: Human evaluation on the VideoJAM-bench dataset. TA, VQ, TC, DD stand for text alignment, visual quality, temporal consistency, and dynamic degree.
@@
C.1 DPO with Human Labels
Human evaluation. Fig. 8 presents additional user study on VideoJAM-bench. Fig. 8(a) shows that StructuralDPO outperforms VanillaDPO in dynamic degree as it performs DPO on video pairs with similar motion. Yet, it underperforms in other dimensions. Fig. 8(b) shows that DenseDPO consistently beats the pre-trained model in all dimensions.
Qualitative results. We show more videos generated by DenseDPO in Fig. 9. Fig. 10 compares DenseDPO with baselines. Overall, DenseDPO aligned model generates videos with high visual quality, rich motion, and precise text alignment. Please check out our project page for video results of baselines and our methods.
C.2 DPO with VLM Labels
@@
Motion bias in VLMs. As discussed in Sec. 3.2, there is a motion bias in human preference annotation—human labelers tend to favor artifact-free slow-motion clips over dynamic clips with artifacts. Ge et al. [21] pointed out that video metrics such as FVD [67] are also biased towards per-frame visual quality rather than temporal motion quality. Here, we study whether more advanced VLM-based video reward models also suffer from this issue. We test two state-of-the-art models, VisionReward [78] and VideoReward [45]. We randomly sample 10k videos from the VanillaDPO training data, each having 121 frames (5s). For each video, we construct a static version of it by duplicating one frame of it, and we compare this static video with the original video.
Tab. 5 presents the winning rate of static vs. original video. Surprisingly, VisionReward favors static videos over original videos with a sizable gap, indicating a clear motion bias. In contrast, VideoReward prefers original videos in around 80% of cases. We note that both VisionReward and VideoReward output multi-dimensional scores (e.g., visual quality, dynamic degree), and aggregate them to predict the binary human preference. VideoReward simply averages all dimensions, and thus is not biased to any dimension. VisionReward instead first labels human preferences on video pairs, and then learns per-dimension weights via logistic regression. This inevitably inherits the motion bias in human labels. Indeed, Tab. 25 in the VisionReward paper [78] reveals that human preference is negatively correlated with object dynamics, while positively correlated with temporal smoothness.
@@
GPT dense preference label. We visualize some dense preferences predicted by GPT o3 Segment in Fig. 11. Overall, GPT is able to identify obvious artifacts such as distorted faces and deformed limbs. With our carefully designed prompt and self-consistency check, it only predicts a preference when there is a clear difference between two segments. However, it still does not understand complex motion, such as playing tennis and cartwheels. This is partially because GPT o3 Segment only has access to 1s video clips, which is too short to finish these actions.
Method VBench Metrics VisionReward Metrics
$s=0.5$ 55.57 57.03 89.88 92.76 92.52 84.25 0.811 0.326 0.601 0.643
$s=1$ (Ours) 55.62 56.13 89.23 92.48 92.99 84.25 0.806 0.305 0.589 0.655
$s=2$ 55.40 56.10 89.04 91.96 92.54 84.15 0.795 0.291 0.557 0.623
C.3 Ablation Study
Dense label granularity. We study the impact of the segment length $s$ in dense preference labels. By default, $s=1$ is used in all our experiments. Here, we tested $s=0.5$ and $s=2$. Due to the high annotation cost, we only label 5k videos for each segment setting in this study. Tab. 6 compares DenseDPO trained on 5k videos using different $s$. $s=1$ consistently outperforms $s=2$ due to more fine-grained preference annotation. Interestingly, $s=0.5$ performs similarly to $s=1$. We hypothesize that this is because $0.5s$ is too short—a longer context window is needed to assess the temporal aspect of videos. In addition, labeling dense preference at $s=0.5$ is 2×2$\times$2 × expensive compared to $s=1$. Therefore, we choose to label 1s segments in our experiments.

### v2

@@ 4.3 DPO with VLM Labels
VLM labelers. As discussed in Sec.˜3.3, we aim to evaluate the effectiveness of pre-trained VLMs in video preference learning. We take VLMs designed specifically for the video quality assessment task, VideoScore [21], LiFT [74], VideoReward [46], and VisionReward [80]. They have been fine-tuned on large-scale human preference labels. We also utilize the state-of-the-art visual reasoning model, GPT o3 [1], to explore the limits of models without task-specific training. Finally, we design GPT o3 Segment that partitions long videos into short segments to process separately, and aggregates results via majority voting. Please refer to Appendix A.6 for more implementation details.
@@
Results. Tab.˜3(a) presents the results of preference prediction. With task-specific fine-tuning, state-of-the-art video reward models outperform the advanced GPT o3 in assessing short clips. Yet, their performance on long videos degrades drastically. Our further analysis in Appendix C.3 reveals that these models might be biased towards video content rather than temporal motion. Thanks to the temporal alignment of video pairs, we can run GPT to compare short segments individually, and then aggregate the results. This leads to higher accuracy in long video preference prediction.
[…]
@@ 4.4 Ablation Study
Human labeling bias. We first study if there is a systematic bias in the annotation pipeline, e.g., labelers may produce higher quality preference labels on short segments than long videos. To verify it, we aggregate labels of all segments within a video to obtain its binary preference label via majority voting, and then train StructuralDPO on these labels. Tab.˜4 shows that it achieves similar results compared to StructuralDPO trained on globally labeled preferences. This proves that DenseDPO’s superior performance comes from segment-level preference supervision instead of labeler bias.
Human label quality. We study how label noise affects DenseDPO performance. As shown in Tab.˜4, we randomly flip 20% and 40% winning or losing labels. This results in a clear drop in all the metrics.
[…]
Ground-truth video in guided generation. In our experiments, we randomly sample 10k videos from the 55k high-quality dataset to serve as guidance videos. Here, we randomly sample another set of 10k videos without overlap with the previously selected ones and conduct DenseDPO training. Tab.˜4 (New GT Video row) shows that both runs achieve similar results, proving that our method is robust to the selection of ground-truth data as long as they are high-quality videos.
Method VBench Metrics VisionReward Metrics
$s=0.5$ 55.57 57.03 89.88 92.76 92.52 84.25 0.811 0.326 0.601 0.643
$s=1$ (Ours) 55.62 56.13 89.23 92.48 92.99 84.25 0.806 0.305 0.589 0.655
$s=2$ 55.40 56.10 89.04 91.96 92.54 84.15 0.795 0.291 0.557 0.623
Dense label granularity. We study the impact of the segment length $s$ in dense preference labels. By default, $s=1$ is used in all our experiments. Here, we tested $s=0.5$ and $s=2$. Due to the high annotation cost, we only label 5k videos for each segment setting in this study. Tab.˜5 compares DenseDPO trained on 5k videos using different $s$. $s=1$ consistently outperforms $s=2$ due to more fine-grained preference annotation. Interestingly, $s=0.5$ performs similarly to $s=1$. We hypothesize that this is because $0.5s$ is too short—a longer context window is needed to assess the temporal aspect of videos. In addition, labeling dense preference at $s=0.5$ is $2\times$ expensive compared to $s=1$. Therefore, we choose to label 1s segments in our experiments.
Table 6: Ablation on DenseDPO training with different amount of GPT o3 segment-level labels. We report VBench [30] and VisionReward [80] metrics on VideoJAM-bench [8].
Method VBench Metrics VisionReward Metrics
10k GPT labels 55.21 56.83 88.31 91.45 92.38 84.20 0.782 0.247 0.360 0.668
35k GPT labels 55.89 58.75 89.68 92.25 92.70 84.89 0.817 0.316 0.498 0.670
55k GPT labels 56.23 60.15 90.75 93.01 92.99 85.21 0.842 0.368 0.598 0.672
VLM label quantity. Finally, we study DenseDPO performance when scaling up automatic VLM labels in Tab.˜6. Since VLM labels are noisy, DenseDPO with 10k GPT labels achieves only marginal improvement compared to the pre-trained model. Nevertheless, as the amount of VLM labels grows, we see consistent improvement in all metrics. This trend reveals that the quantity of VLM labels compensates for their quality, and our method scales well with the quantity of automatic labels. Since VLM itself is an actively evolving field, we view this as a promising future direction to further scale up automatic preference learning with more data and better off-the-shelf VLMs.
@@ 5 Conclusion
We present DenseDPO, an improved preference optimization framework for video generation. We address two critical aspects of video DPO—comparison data curation and preference labeling. Our guided video generation mechanism and fine-grained preference labeling significantly improve over vanilla DPO. Furthermore, we show that segment-level labels unlock automatic annotation with off-the-shelf VLMs without task-specific training. We discuss our limitations in Appendix D.
@@ A.2 Baselines
[…]
@@
REFL [79] directly maximizes the reward from a reward model via gradient ascent. We tried REFL in our preliminary experiments with HPSv2 [75] and VisionReward [80] as the target reward models. However, we observed severe reward hacking, e.g., the video frames become oversaturated and the temporal motion gets static. It also requires huge GPU memory and computation as we need to decode latent tokens to raw pixels to apply these VLMs and backpropagate gradients through them. Since the results are clearly worse than DPO, we did not continue this direction and do not report their results in the paper.
SFT fine-tunes the pre-trained model on the 55k high-quality dataset described in Sec.˜A.1 for 5k iterations. We did not observe a clear difference between full-model and LoRA [27] fine-tuning, and thus choose LoRA fine-tuning as it is more efficient.
D3PO [83] shares a similar objective function as VanillaDPO. We adapt their official codebase [^1] https://github.com/yk7333/d3po to our video setting, and use the same preference labels as in VanillaDPO.
DPOK [17] is an online RL-based method. We also adapt their official codebase [^2] https://github.com/google-research/google-research/tree/master/dpok to our video setting, and leverage VisionReward [80] to provide online reward feedback. In our experiments, we found that DPOK is unstable to train and easily results in reward hacking. Thus, we use a low learning rate of $2\times 10^{-6}$ and a high KL weight to prevent training divergence.
SePPO [90] can be viewed as a combination of DPO and SFT, as its preference data come from samples generated by a past checkpoint of the current model and pre-collected high-quality data. Since no training code is available for SePPO, we re-implement it based on its paper.
@@
DPO training hyper-parameters. Both VanillaDPO and StructuralDPO adopt the Flow-DPO loss with a constant $\beta$, as prior work shows it achieves the best performance [46]. Following prior works [80, 46], we use $\beta=500$ and apply LoRA [27] with rank 128 to fine-tune the video model. We train with the AdamW optimizer [49] and a global batch size of 256 for 1k steps. The peak learning rate is set to $1\times 10^{-5}$ and it is linearly warmed up from 0 in the first 250 steps. A gradient clipping of 1.0 is applied to stabilize training. We implement all models using PyTorch [54] and conduct training on 64 NVIDIA A100 GPUs, which takes around 16 hours.
Inference hyper-parameters. We use the rectified flow sampler [48] with 40 sampling steps and a classifier-free guidance [22] scale of 88 to generate horizontal videos of $512\times 288$ resolution. A timestep shifting [16] of $5.66$ is applied to further improve results.
@@ A.4 Evaluation Datasets
At the beginning of the project, we experimented with the prompt suite of VBench [30]. However, we noticed that the model can achieve high scores with good visual quality and even degraded motion dynamics. Since our main focus is improving the temporal quality of the pre-trained video model and VanillaDPO, we choose to evaluate on two motion-rich prompt sets.
@@
MotionBench collects more diverse prompts from existing prompt sets [58, 35, 74]. We run GPT-4o to select prompts with dynamic actions as described in Sec.˜A.1, resulting in 419 prompts. This prompt set contains more diverse scenes, subjects, and action types.
@@ B.1 Analyzing the Learning Signal
As discussed in Sec.˜3.2, guided sampling is attractive since it fixes the structure in the preference pair. This neutralizes the motion bias between videos and focuses the comparison on visual artifacts. However, StructuralDPO with guided sampling achieves inferior results. Analyzing its learning signal reveals that it can paradoxically push the model to “unlearn” the real data distribution. Intuitively, this happens because the learning signal from a losing sample typically dominates over the winning one in those regions of a video, which correspond to the real data distribution. In this section, we investigate this phenomenon and discuss potential remedies to the issue.
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},\,t\sim\mathcal{U}(0,T),\,\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}\mid\bm{x}_{0}^{w}),\,\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}\mid\bm{x}_{0}^{l})}\Big[\log\sigma\Big(-\beta T\big($ (8)
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},\,t\sim\mathcal{U}(0,T),\,\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}\mid\bm{x}_{0}^{w}),\,\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}\mid\bm{x}_{0}^{l})}\Big[\log\sigma\Big(-\beta T\omega(t)\big($ (11)
@@
In guided sampling (see Algo.˜2), we generate a winning/losing sample $\bm{x}^{*}_{n}=(1-\eta)\bm{x}_{0}+\eta\bm{\epsilon}^{*}$ using a real video $\bm{x}_{0}$. This makes them carry similar structure, which means that the winning and losing samples share many pixels. Moreover, these shared pixels normally correspond to the ground-truth data distribution. Let’s split the pixels into two sets, $\mathcal{I}_{\text{same}}$ and $\mathcal{I}_{\text{unique}}$:
$\mathcal{I}_{\text{same}}(\bm{x}^{w},\bm{x}^{l})$ $=\{p\mid\bm{x}^{w}_{0}[p]\approx\bm{x}^{l}_{0}[p]\approx\bm{x}_{0}[p]\},$ (18)
$\mathcal{I}_{\text{unique}}(\bm{x}^{w}_{0},\bm{x}^{l}_{0})$ $=\{p\mid p\not\in\mathcal{I}_{\text{same}}\},$ (19)
where $p$ is a pixel location (e.g., typically, triplet $(i,j,k)$ denoting the frame, height, width indices). In this way, $\mathcal{I}_{\text{same}}$ stores pixel locations which remained intact during the forward diffusion process of guided sampling and subsequent denoising with $\mathbf{G}_{\bm{\theta}}$. Fig.˜2 and Fig.˜11 illustrate this: many pixels in the winning and losing samples are identical and correspond to the original ground-truth video. In this way, one can argue that $\mathcal{I}_{\text{same}}$ corresponds to the real data distribution.
@@ B.2 Empirical Verification
The resulting dataset comprises five image variants, progressively more corrupted, visualized in Fig.˜6 (top). This setup allows separate measurement of losses in corrupted and uncorrupted regions. Subsequently, we randomly sample a timestep $t\sim\mathcal{U}[0,1]$, estimate velocities $\bm{v}_{\text{Flux}}(\bm{x}_{t}^{w},t)$ and $\bm{v}_{\text{Flux}}(\bm{x}_{t}^{l},t)$, and compute per-pixel losses (averaged over the 16 latent channels $C$):
@@
Loss computations use the same noise seed for the corrupted and uncorrupted images (as is usually done in practice). Fig.˜6 (top) illustrates a representative example of corruption. Visualizations in Fig.˜6 (bottom) clearly demonstrate the loss dominance of the losing samples in the uncorrupted regions. As we maximize the loss of the losing sample and minimize that of the winning sample, this results in the model increasing loss in uncorrupted image regions when trained with DPO to mitigate blurring artifacts.
Fig.˜7 provides a quantitative analysis, demonstrating how increasing artifact severity intensifies the negative learning signal from losing samples in uncorrupted image areas.
@@ Appendix C Additional Experimental Results
Table 7: Quantitative comparison with more baselines on VideoJAM-bench [8]. We report automatic metrics from VBench [30] and VisionReward [80].
Method VBench Metrics VisionReward Metrics
D3PO [83] 56.15 58.03 88.93 92.23 92.78 82.53 0.833 0.322 0.482 0.602
DPOK [17] 54.99 56.28 89.14 92.36 92.95 78.65 0.795 0.337 0.518 0.457
SePPO [90] 55.83 57.42 89.93 92.65 92.93 82.85 0.841 0.326 0.554 0.587
DenseDPO 56.99 60.92 91.54 93.84 93.56 85.38 0.863 0.376 0.632 0.680
@@
Figure 8: Human evaluation on the VideoJAM-bench dataset. TA, VQ, TC, $D$ stand for text alignment, visual quality, temporal consistency, and dynamic degree.
@@
C.1 Comparison with Additional Baselines
In Tab.˜7, we compare with additional baselines on VideoJAM-bench [8]. D3PO [83] achieves similar results as VanillaDPO since they are both DPO-based methods. DPOK [17], an online RL method, slightly improves the visual quality and temporal consistency compared to the pre-trained model, yet significantly degrades the dynamic degree. This is because the reward model it uses, VisionReward [80], is biased towards per-frame visual quality instead of temporal motion (see more analysis in C.3). Thus, training with such online feedback leads to static motion. SePPO [90] achieves a better dynamic degree than VanillaDPO, while performing worse in visual quality. Instead of using offline-generated losing samples, SePPO generates online samples with a reference model, and designs a filtering method (AAF) to assess its quality. However, we noticed that AFF is based on model denoising loss, which is sometimes unreliable. In those cases, the model is optimized on low-quality videos, leading to degraded visual quality.
C.2 DPO with Human Labels
Human evaluation. Fig.˜8 presents an additional user study on VideoJAM-bench. Fig.˜8(a) shows that StructuralDPO outperforms VanillaDPO in dynamic degree as it performs DPO on video pairs with similar motion. Yet, it underperforms in other dimensions. Fig.˜8(b) shows that DenseDPO consistently beats the pre-trained model in all dimensions.
Qualitative results. We show more videos generated by DenseDPO in Fig.˜9. Fig.˜10 compares DenseDPO with baselines. Overall, DenseDPO aligned model generates videos with high visual quality, rich motion, and precise text alignment. Please check out our project page for video results of baselines and our methods.
C.3 DPO with VLM Labels
@@
Motion bias in VLMs. As discussed in Sec.˜3.2, there is a motion bias in human preference annotation—human labelers tend to favor artifact-free slow-motion clips over dynamic clips with artifacts. Ge et al. [20] pointed out that video metrics such as FVD [69] are also biased towards per-frame visual quality rather than temporal motion quality. Here, we study whether more advanced VLM-based video reward models also suffer from this issue. We test two state-of-the-art models, VisionReward [80] and VideoReward [46]. We randomly sample 10k videos from the VanillaDPO training data, each having 121 frames (5s). For each video, we construct a static version of it by duplicating one frame of it, and we compare this static video with the original video.
Tab.˜8 presents the winning rate of static vs. original video. Surprisingly, VisionReward favors static videos over original videos with a sizable gap, indicating a clear motion bias. In contrast, VideoReward prefers original videos in around 80% of cases. We note that both VisionReward and VideoReward output multi-dimensional scores (e.g., visual quality, dynamic degree), and aggregate them to predict the binary human preference. VideoReward simply averages all dimensions, and thus is not biased to any dimension. VisionReward instead first labels human preferences on video pairs, and then learns per-dimension weights via logistic regression. This inevitably inherits the motion bias in human labels. Indeed, Tab. 25 in the VisionReward paper [80] reveals that human preference is negatively correlated with object dynamics, while positively correlated with temporal smoothness.
@@
GPT dense preference label. We visualize some dense preferences predicted by GPT o3 Segment in Fig.˜11. Overall, GPT is able to identify obvious artifacts such as distorted faces and deformed limbs. With our carefully designed prompt and self-consistency check, it only predicts a preference when there is a clear difference between two segments. However, it still does not understand complex motion, such as playing tennis and cartwheels. This is partially because GPT o3 Segment only has access to 1s video clips, which is too short to finish these actions.
C.4 Failed Attempts
Here, we record some failed approaches we tried.
Real videos as winning data. Since our goal is to improve the motion quality of video models, real-world videos can naturally serve as winning samples as they follow physical laws perfectly. We tried DPO using our 55k high-quality training videos as winning data, and videos generated by our model as losing data. The implicit accuracy quickly converges to almost 100% within 500 steps, yet the model generation does not improve. This is likely because real videos are significantly better than the generated ones, which fails to provide useful signals to improve the video model.
Video pairs with different guidance $\eta$. Intuitively, videos generated with more guidance are often better than those with lower guidance. We tried generating video pairs with different $\eta$ and assign winning-losing labels based on it. This gives us preference labels “for-free". However, DPO on this data does not improve model generation. Our hypothesis is that the model may infer the noise level from the generated samples instead of measuring its quality.